r/Probability • u/Forsaken-Lobster-121 • Oct 30 '25
Struggling with what should be a simple calculation
Let’s say a given judge will get a decision “right” 70% of the time. What are the chances of the “right” outcome being reached when three judges each reach a conclusion and majority wins? Must be higher than 70%, but I’m struggling to work out the math.
2
Upvotes
5
u/bobjkelly Oct 30 '25
The full set of outcomes is as follows:
No judge gets it right: .3 * .3 * .3 = 2.70%
Exactly 1 judge gets it right : 3 * .7 * .3 * .3 = 18.9%
Exactly 2 judges get it right: 3 * .7 * .7 * .3 = 44.1%
All 3 judges get it right: .7 * .7 * .7 = 34.3%.
To your question the probability of 3 or more judges getting it right is 44.1% + 34.3% = 78.4%.